Find the equation of the circle with endpoints at (1,3) and (-5, -5)
Answer by stanbon(75887)

Find the equation of the circle with endpoints at (1,3) and (-5, -5)
---
Center = midpoint = (-2,-1)
radius = sqrt[(1--5)^2+(3--5)^2] = sqrt[36+64] = 10
---
Equation:
(x+2)^2 + (y+1)^2 = 100
============================
Cheers,
Stan H.